# PromptKD: Unsupervised Prompt Distillation for Vision-Language Models

### v4

@@ 1 Introduction
Recently large pretrained vision-language models (VLMs), such as CLIP [37] and ALIGN [16], have demonstrated superior generalization ability for domain-specific downstream tasks. Unlike conventional visual frameworks, the vision-language model, like CLIP, usually employs a two-tower architecture that includes an image encoder and a text encoder. These models are trained using a contrastive loss to learn a unified embedding space that aligns the representations of multi-modal signals.
@@ 2 Related Work
Prompt Learning in Vision-Language Models. Prompt learning is a technique that can transfer the large pre-trained model, like CLIP [37], towards downstream tasks [61, 38, 10] without the need for completely re-training the original model. It proposes to adapt the representations for specific tasks through learnable text or visual soft prompts instead of manually crafted hard prompts (e.g., “a photo of a {classname}”). Soft prompts [17, 66, 65, 23, 40] can be optimized by back-propagating through the frozen pre-trained model, resulting in better performance. Existing works mainly focus on designing various efficient forms of prompts using scarce labeled domain data. MaPLe [19] proposes to learn prompts for the image and text branches simultaneously, rather than a separate side. PromptSRC [20] utilizes its original features to regularize the learning of prompts for each branch. Previous works necessitated forward and backward computations for each input in both image and text branches. In our work, we leverage the unique decoupled-modality characteristic of CLIP, saving well-trained teacher text features as class vectors for student distillation. In this way, the training of student CLIP is simplified to solely include forward and backward calculations of the image branch, without requiring the text branch.
@@
Zero-shot Learning. Given the labeled training set of the seen classes, zero-shot learning (ZSL) [53, 51] aims to learn a classifier that can classify testing samples of unseen classes. Existing methods can be roughly divided into two types based on whether test images are available: Inductive [62, 54] and Transductive [45, 47] ZSL. Previous works on prompt learning, such as MaPLe and PromptSRC, have mainly focused on the instance inductive settings where only labeled training instances are available. In our paper, we explore the transductive ZSL setting where both seen and unseen class images are all utilized in model learning. Specifically, our teacher model follows the same training scheme as PromptSRC, which is trained on samples from seen classes with ground truth labels. The difference is that the target student model is trained on the full unlabeled dataset, which contains all samples of both seen and unseen classes, without using any ground truth labels.
@@
Knowledge Distillation. Knowledge distillation [14] aims to train a lightweight student model under the supervision of a large pretrained teacher model. In recent years, various distillation forms have emerged for effective knowledge transfer from teachers to students, such as logits alignment [63, 26, 64, 28], feature imitation [59, 4] and sample relationship matching [34, 56]. In addition to traditional image classification topics, knowledge distillation has achieved great success in many vision tasks, including object detection [2, 50], image segmentation [29, 57], and pose estimation [27]. Recently, many works [36, 58, 52, 22] have turned their attention to the CLIP model. These works propose leveraging the CLIP model’s exceptional generalization capabilities to enhance the learning of existing models. CLIP-KD [58] find that in distilling pre-trained CLIP models, the simplest feature mimicry with the MSE loss approach yields the best results. TinyCLIP [52] performs cross-modal feature alignment in affinity space between teacher and student. Our approach differs from previous distillation methods that train the entire student model by leveraging a pre-trained large CLIP teacher. In our work, we employ a more efficient approach by utilizing student prompts for distillation while keeping the student’s original CLIP weights frozen. This allows us to achieve the desired knowledge transfer without the need for extensive re-training of the student model.
@@ 7 Additional Experiments
StandfordCars 84.53
Flowers102 99.05
Food101 94.56
FGVCAircraft 54.44
SUN397 84.97
DTD 85.76
EuroSAT 94.79
UCF101 89.50

### v5

@@ 1 Introduction
Recently large pretrained vision-language models (VLMs), such as CLIP [41, 68] and ALIGN [17], have demonstrated superior generalization ability for domain-specific downstream tasks. Unlike conventional visual frameworks, the vision-language model, like CLIP, usually employs a two-tower architecture that includes an image encoder and a text encoder. These models are trained using a contrastive loss to learn a unified embedding space that aligns the representations of multi-modal signals.
@@ 2 Related Work
Prompt Learning in Vision-Language Models. Prompt learning is a technique that can transfer the large pre-trained model, like CLIP [41], towards downstream tasks [66, 42, 11] without the need for completely re-training the original model. It proposes to adapt the representations for specific tasks through learnable text or visual soft prompts instead of manually crafted hard prompts (e.g., “a photo of a {classname}”). Soft prompts [18, 72, 71, 25, 44] can be optimized by back-propagating through the frozen pre-trained model, resulting in better performance. Existing works mainly focus on designing various efficient forms of prompts using scarce labeled domain data. MaPLe [21] proposes to learn prompts for the image and text branches simultaneously, rather than a separate side. PromptSRC [22] utilizes its original features to regularize the learning of prompts for each branch. Previous works necessitated forward and backward computations for each input in both image [8, 56] and text branches. In our work, we leverage the unique decoupled-modality characteristic of CLIP, saving well-trained teacher text features as class vectors for student distillation. In this way, the training of student CLIP is simplified to solely include forward and backward calculations of the image branch, without requiring the text branch.
@@
Zero-shot Learning. Given the labeled training set of the seen classes, zero-shot learning (ZSL) [58, 55, 32] aims to learn a classifier that can classify testing samples of unseen classes. Existing methods can be roughly divided into two types based on whether test images are available: Inductive [67, 59] and Transductive [49, 51] ZSL. Previous works on prompt learning, such as MaPLe and PromptSRC, have mainly focused on the instance inductive settings where only labeled training instances are available. In our paper, we explore the transductive ZSL setting where both seen and unseen class images are all utilized in model learning. Specifically, our teacher model follows the same training scheme as PromptSRC, which is trained on samples from seen classes with ground truth labels. The difference is that the target student model is trained on the full unlabeled dataset, which contains all samples of both seen and unseen classes, without using any ground truth labels.
@@
Knowledge Distillation. Knowledge distillation [15] aims to train a lightweight student model under the supervision of a large pretrained teacher model. In recent years, various distillation forms have emerged for effective knowledge transfer from teachers to students, such as logits alignment [69, 29, 70, 31], feature imitation [64, 4, 27] and sample relationship matching [38, 61]. In addition to traditional image classification topics, knowledge distillation has achieved great success in many vision tasks, including object detection [2, 54, 19], image segmentation [33, 62], and pose estimation [30]. Recently, many works [40, 63, 57, 24] have turned their attention to the CLIP model. These works propose leveraging the CLIP model’s exceptional generalization capabilities to enhance the learning of existing models. CLIP-KD [63] find that in distilling pre-trained CLIP models, the simplest feature mimicry with the MSE loss approach yields the best results. TinyCLIP [57] performs cross-modal feature alignment in affinity space between teacher and student. Our approach differs from previous distillation methods that train the entire student model by leveraging a pre-trained large CLIP teacher. In our work, we employ a more efficient approach by utilizing student prompts for distillation while keeping the student’s original CLIP weights frozen. This allows us to achieve the desired knowledge transfer without the need for extensive re-training of the student model.
@@ 7 Additional Experiments
StandfordCars 77.70
Flowers102 77.54
Food101 91.59
FGVCAircraft 31.29
SUN397 70.86
DTD 56.32
EuroSAT 47.55
UCF101 76.20
Avg. 71.89